OpenStudy (anonymous):

Find the center of the circle with equation x^2+y^2-10x+6y+27=0

sam (.sam.):

You gotta put it in this form \[(x-h)^2+(y-k)^2=r^2\] Try using completing the square \[ [x^2-10x]+[y^2+6y]+27=0\]

sam (.sam.):

Then h and k is the center